The pulse rates of 147 randomly selected adult males vary from a low of 45 bpm to a high of 105 bpm. Find the minimum sample size required to estimate the mean pulse rate of adult males. Assume that we want 95% confidence that the sample mean is within 4 bpm of the population mean. Complete parts (a) through (c) below. a. Find the sample size using the range rule of thumb to estimate (Round up to the nearest whole number as needed.) b. Assume that ơ 11.9 bpm, based on the value s-1 1 .9 bpm from the sample of 147 male pulse rates. (Round up to the nearest whole number as needed.) c. Compare the results from parts (a) and (b). Which result is likely to be better? The result from part (a) is the result from part (b). The result fromis likely to be better because the same size as arger than smaller than
Answer #1

a) Estimated SD = (105 - 45)/4 = 15

E = 4

For 95% confidence, z = 1.96

Hence,

Sample size required

1.96 152

n = 55

b) SD = 11.9

E = 4

For 95% confidence, z = 1.96

Hence,

Sample size required

n = (\frac{1.96*11.9}{4})^2

n = 35

c) Larger than; part b); it gives a more reasonable estimate.